If Gama is the ratio of educate of an ideal gas the degree the number of degree of freedom of molecule of the gases

Grade:11

1 Answers

Mohammed Kasib Siddiqui
5 years ago
Gamma= 1+ 2/f      where f is the no. Of dehree of                                       freedom
Hence
Gamma-1= 2/f 
(Gamma-1)/2= 1/f
f= 2/(gamma-1)